Given: f(x)=-2+5 Find:
a. f(0)
b. f(2)
c. f(-3)
d. f(y)

Answer by funmath(2933)

I think you must have meant to put an x in after the -2 otherwise the answers would all be 3.
Given: f(x)=-2x+5 Find:
a. f(0)=-2(0)+5
=0+5
=5
b. f(2)=-2(2)+5
=-4+5
=1
c. f(-3)=-2(-3)+5
=6+5
=11
d. f(y)=-2(y)+5
=-2y+5
Happy Calculating!!!